You are appointed as manager of a wood panel cutting shop in a furniture manufacturing plant. You are responsible for implementing a TPM approach. and we are presenting the following data (defined from the last 12 months). - The workshop is shut down for an average of 50 hours per month for 350 scheduled hours of operation. The shutdowns are distributed as follows : • 15 hours for production changeover. • 20 hours for machine stop. • 10hours for quality problems. • 5hours for unknown origine - The workshop produces an average of 9000 pieces per month instead of 34 pieces per hour. - The quality rate is 93%. a) Calculate the workshop OEE and each rate. b) What actions will you launch as a priority and why? c) What difficulties do you think you are encountering in implementing these actions? d) How to overcome the difficulties?
Answer #1

OEE = availability x quality x efficiency

Here, availability = 300/350 = 0.85714

quality = 0.93

Efficiency = pieces actually produced / theoritical production

= 9000 / 34 x300 = 0.8823

OEE = 0.88235 x 0.85714 x0.93 = 0.7033

(b) The biggest contributor to the low OEE is the quality related issues, which need to be corrected at many levels

The downtime can be reduced by employing techniques like SMED, which reduce the changeover time of production. The other way is to employ preventive maintenance of the machines so as to minimise the stops, and training the workforce to repair the machines in shorter time. Another way is to solve the quality related problems by implementing a comprehensive quality system and reduce the stoppage time. The remaining part of downtime can be studied further t find its origin and act accordingly. Among other actions will be the staff training and making the quality first concern of each employee, by checking at multiple stages. Resolving the quality issue will prove majority of problems.

(c) The possible problems in implementing the changes will be the inetrtia from the employees who would be susceptible due to a number of reasons, financial implications of the change and management's willingness to go ahead with change.

(d) The inertia to change on the part of employees can be overcome by educating and traaining them about the changes and the positive impacts of those changes on company and its employees, while management's position can be change can be influenced by providing them detailed and rational analysis of facts and figures and the expected benefits from the changes.